# Cryogenic Thermal Modeling of Microwave High Density Signaling

### v1

@@ IV.4 Method of Calculating Active Heat Load
IV.5 Active Load Example in a Bluefors XLD1000-SL Dilution Refrigerator
For this example, we will take a typical attenuation configuration with 20 dB attenuation at the 4K plate of the fridge Krinner et al. (2019) and the line lengths from a Bluefors XLD1000-SL system detailed in Table 1. Using the experimentally determined cable resistivity and a flux bias current of I = 0.4 mA, the calculated active heat loads of a single SC-086/50-SCN-CN cable in the Bluefors XLD1000-SL system are shown in Table 9.
@@ IV.5 Active Load Example in a Bluefors XLD1000-SL Dilution Refrigerator
Figure 7: (a) Diagram showing the layout of the model systems considered in this thermal model with a square array of $n\times n$ flux tunable transmons and a tunable coupler between each pair of transmons. Each model has $n^{2}$ transmons and $2n\left(n-1\right)$ tunable couplers. (b) The fridge wiring diagram of the model system shows the assumed attenuator configuration for each type of line at each fridge stage and the readout chain electronics, including TWPA and LNA amplifiers at the MXC and 4K stages, respectively. Note that the attenuators in the Read-In and Qubit XY lines are just shown without values because they are irrelevant to the thermal model.
@@ 15 $\times$ 15 225 420 38 450 420 114 984
Table 10: The sizes of the different model quantum processors considered. Each processor is assumed to be an $n\times n$ square array containing $n^{2}$ qubits, with tunable couplers between each pair of qubits. Each qubit requires two control lines: the XY drive and flux bias lines, while the tunable couplers require a single flux bias line each. The readout circuit requires three lines: a read-in line, a read-out line, and one line for the TWPA pump, and is assumed to be six-way multiplexed. The next entry in the table would be a 16 $\times$ 16 array with 256 qubits, but this would require a total of 1121 lines, which is beyond the capacity of $N$ = 1008 lines available in the dilution fridge. Each thermal model assumes that only these lines are present and there are no unused lines in the system.
@@ V Application of the Thermal Model to a Bluefors XLD1000-SL Dilution Refrigerator
This section estimates the full thermal load of different-sized quantum computers with between 100 and 225 qubits installed in a Bluefors XLD1000-SL dilution refrigerator. It does not use the full number of cables the fridge can support. Rather, it estimates real systems using only those signal cables needed that the available cooling power could accommodate.
The model quantum processors are shown in Fig. 7(a) and consist of a square array of $n\times n$ flux tunable transmon qubits, where $n=10,11,...,15$, with tunable couplers between each pair Yan et al. (2018). Each model quantum processor has $n^{2}$ qubits, and $2n\left(n-1\right)$ tunable couplers. Each transmon is assumed to require two control lines - a flux bias line and an XY drive line - and is also coupled to its own output resonator. Each tunable coupler has its own flux bias line. Model processors with an array of up to 15 $\times$ 15 = 225 qubits can be accommodated in the fridge, since a 16 $\times$ 16 array with 256 qubits would require a total of 1121 lines, which is beyond the capacity of $N$ = 1008 lines available.
Figure 8: (a) The fraction of the estimated cooling powers at the different fridge stages used by the total heat loads for model quantum processor configurations of various sizes. (b) An example of the fraction of the estimated cooling powers used at each stage for a 144 qubit processor.
Each readout circuit has a read-in line to the on-chip resonators that continues to a readout line with isolators, TWPA O’Brien et al. (2014); Macklin et al. (2015); Feng et al. (2020) fed by a pump line coming into a directional coupler, and a semiconductor LNA at the 4K stage. The readout chain assumes six-way multiplexing of the readout resonators onto $ceil\left(n^{2}/6\right)$ separate readout circuits, where $ceil$ is the ceiling function, which rounds a fraction up to the next integer. Each of the readout circuits requires three microwave lines: one read-in, one read-out, and one line for the TWPA pump. When $n^{2}$ is not a multiple of six, one readout circuit will have fewer than six qubits on it. Real designs of processors usually pick the number of qubits to be a multiple of the readout multiplexing to simplify routing, whereas these model processors are just selected to illustrate the scaling of the thermal load. Fig. 7(b) shows the fridge wiring diagram of the model system with the assumed attenuator configuration for each type of line at each fridge stage and the specifics of the readout chain electronics, including TWPA and LNA amplifiers at the MXC and 4K stages, respectively.
@@
In Fig. 8(a) we show the total heat loads as a fraction of the available cooling power at each stage as a function of processor size, using the estimated cooling powers at each stage of a Bluefors XLD1000-SL dilution refrigerator shown in Table 1. Fig. 8(b) shows a typical simulation result for 144 qubits, where the origin of the heat loads at each stage is shown in a stacked bar graph. In principle, the system will be able to cool to base temperature provided all stages have less heat load than their nominal cooling power.
@@
We caution readers that the model does not capture some additional heat loads that will change the results. In particular, the active heat load due to the resistance of the package that holds the processor is not considered here. Other heat loads, such as the presence of a long-life cold trap, can add further static loads to the 50K and 4K stages. As such, this model only captures a general overview, and we use it to demonstrate a system with sufficient engineering margin that could be used as a starting point for a real design. The full configuration and sources of heat would need to be modeled before the authors would have sufficient confidence to recommend any particular design. In addition, the time taken to cool to base temperature may be an important system parameter. This thermal model only considers the steady state case and a more complete model would be needed to determine the cooldown time, which will depend on the mass, heat capacity, and thermal conductances of each item attached to a stage.
@@
Options to improve the thermal performance include replacing the SC-086/50-SCN-CN cable with microwave flex cable that reduces the static load by scaling the cross-sectional area and combining several signals into one assembly or with cables made from other low thermal conductivity metals such as brass or stainless steel. In this case, there is a trade-off between reducing the static load due to lower thermal conductivity and increasing the active load due to higher electrical resistivity. Using a superconducting cable would dramatically reduce the static and active heat loads, particularly between the 4K and the MXC stages. This option is possible at the lower fridge stages below the 4K stage, where the temperatures are below the superconducting transition temperatures of commercially available superconducting cables zot (b). The TWPA pump termination resistor could also be relocated to the cold plate, which has a lower heat load. New designs of kinetic inductance parametric amplifiers operate at 4 K and may be able to replace the current LNA and the TWPA Malnou et al. (2021). Optical transduction of the microwave output signal at the mixing chamber van Thiel et al. (2023) is another method that would allow the LNA amplifiers to be removed and replace the microwave readout chain with an optical one. Finally, control lines using optical fibers and transducers to create RF signals at the base of the fridge have also been demonstrated and could potentially replace the coax cables Lecocq et al. (2021).
@@
In addition, the required volume, mass, and mounting infrastructure for typical off-the-shelf readout components such as circulators and TWPA amplifiers could potentially put an upper bound on the number of readout chains that could fit in the MXC, which in turn will put an upper bound on the number of qubits possible for such system. Based on experience with the current hardware size, it would be challenging to accommodate more than the 24 readout lines required for the 144-qubit processor on the mixing chamber of a Bluefors XLD1000-SL system.
@@ VI Conclusion
We measured the thermal conductivity and electrical resistivity of SC-086/50-SCN-CN coaxial cables from room temperature to approximately 4 K. This data was then used to create smooth functions of each respective property with temperature, using simple approximations to extend the fits beyond the experimental data down to the millikelvin regime. We then used these functions to calculate static and active heat loads for a superconducting quantum processor in a dilution fridge in realistic configurations. Finally, we applied these thermal models to a set of hypothetical model flux tunable superconducting transmon processors of increasing size. The model processors use the tunable coupler architecture, and so each transmon requires two signal lines (an XY control and a flux bias line), while the tunable couplers require a flux bias line. The readout system is also modeled, including TWPA amplifiers at the mixing chamber and semiconductor LNA amplifiers at the 4K stage. For a Bluefors XLD1000-SL dilution refrigerator using HDW wiring, the theoretical limit of the system using all the available cooling power is approximately 200 qubits. However, with engineering margin in the cooling power and the available space for microwave readout circuitry at the mixing chamber, the practical limit is approximately 140 qubits.
VII Acknowledgements
We would like to thank Jen-Hao Yeh, Sanya Deshpande, and Michael Selvanayagam for their guidance in the early stages of this project.
This work was produced by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy. Publisher acknowledges the U.S. Government license to provide I-238 public access under the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).

### v3

@@ IV.3 Method of Calculating Active Heat Load
IV.4 Active Load Example in a Bluefors XLD1000-SL Dilution Refrigerator
For this example, we take a typical attenuation configuration with 20 dB attenuation at the 4K plate of the fridge krinner2019 and the line lengths from a Bluefors XLD1000-SL system detailed in Table 1. Using the experimentally determined cable resistivity and a flux bias current of I = 0.4 mA, the calculated active heat loads of a single SC-086/50-SCN-CN cable in the Bluefors XLD1000-SL system are shown in Table 9.
@@ IV.4 Active Load Example in a Bluefors XLD1000-SL Dilution Refrigerator
Figure 7: (a) Diagram of the layout of the model systems considered in this thermal model, which consists of a square array of $n\times n$ flux-tunable transmons and a tunable coupler between each pair of transmons. Each model has $n^{2}$ transmons and $2n\left(n-1\right)$ tunable couplers. (b) The fridge wiring diagram of the model system, in which the assumed attenuator configuration for each type of line at each fridge stage and the readout chain electronics, including TWPAs and LNAs at the MXC and 4K stages, respectively, are shown. Note that the attenuators in the Read-In and Qubit XY lines are shown without values because they are irrelevant to the thermal model.
@@ 15 $\times$ 15 225 420 38 450 420 114 984
Table 10: The sizes of the different model quantum processors considered. We assume each processor to be an $n\times n$ square array containing $n^{2}$ qubits, with tunable couplers between each pair of qubits. Each qubit requires two control lines: the XY drive and flux bias lines, while the tunable couplers require a single flux bias line each. The readout circuit requires three lines - a read-in line, a read-out line, and one line for the TWPA pump - and we assume that it is six-way multiplexed. Finally, each thermal model assumes that only these lines are present and there are no unused lines in the system.
@@ V Application of the Thermal Model to a Bluefors XLD1000-SL Dilution Refrigerator
In this section, we estimate the full thermal load of different-sized quantum computers with between 100 and 225 qubits installed in a Bluefors XLD1000-SL dilution refrigerator. For this estimate, we do not use the full number of cables the fridge can support. Instead, we calculate estimates for real systems using only those signaling cables required to control a particular size of processor.
The model quantum processors are shown in Fig. 7(a) and consist of a square array of $n\times n$ flux tunable transmon qubits, where n=10,11,,15n=10,11,...,15, with tunable couplers between each pair yan2018. Each model quantum processor has $n^{2}$ qubits, and $2n\left(n-1\right)$ tunable couplers. We assume that each transmon requires two control lines - a flux bias line and an XY drive line - and that they are also each coupled to their own output resonator. Each tunable coupler has its own flux bias line. Model processors with an array of up to 15 $\times$ 15 = 225 qubits can be accommodated in the fridge, since a 16 $\times$ 16 array with 256 qubits would require a total of 1121 lines, which is beyond the capacity of $N$ = 1008 lines available.
Figure 8: (a) The fraction of the estimated cooling powers at the different fridge stages used by the total heat loads for model quantum processor configurations of various sizes. (b) An example of the fraction of the estimated cooling powers used at each stage for a 144-qubit processor. The LNA heatload only appears at the 4K stage, where it is situated. There are TWPA loads at the MXC due to the 50 Ohm termination resistor on the TWPA pump line, and at the 4K, Still, and CP stages due to attenuators in the TWPA pump line at those stages. The TWPA loads at the 50K and 4K stages are small, even at the 4K stage with an attenuator, and cannot be easily seen in the bar chart. Similarly, the active loads at the 50K, Still, and CP are very small fractions and cannot be easily seen in the bar chart. There are no attenuators in the flux bias lines at those stages, and the active load is due only to the cable resistance.
Each readout circuit has a read-in line to the on-chip resonators that continues to a readout line with isolators, TWPA obrien2014; macklin2015; feng2020 fed by a pump line coming into a directional coupler, and a semiconductor LNA at the 4K stage. The readout chain assumes six-way multiplexing of the readout resonators onto $ceil\left(n^{2}/6\right)$ separate readout circuits, where ceilceil is the ceiling function, which rounds a fraction up to the next integer. Each of the readout circuits requires three microwave lines: one read-in, one read-out, and one line for the TWPA pump. When $n^{2}$ is not a multiple of six, one readout circuit has fewer than six qubits on it. Real designs of processors usually pick the number of qubits to be a multiple of the readout multiplexing to simplify routing, whereas these model processors are just selected to illustrate the scaling of the thermal load. The fridge wiring diagram of the model system with the assumed attenuator configuration for each type of line at each fridge stage and the specifics of the readout chain electronics, including TWPAs and LNAs at the MXC and 4K stages, respectively, are shown in Fig. 7(b).
@@
In Fig. 8(a), we show how the total heat loads as a fraction of the available cooling power at each stage vary with increasing processor size, using the estimated cooling powers at each stage of a Bluefors XLD1000-SL dilution refrigerator shown in Table 1. Fig. 8(b) presents a typical simulation result for 144 qubits, where the origin of the heat loads at each stage is shown in a stacked bar graph. In principle, the system has the ability to cool to base temperature provided all stages have less heat load than their nominal cooling power.
Surprisingly, the largest fractional heat load appears at the 4K stage. In most dilution fridge experiments, the heat load at the MXC stage is the most critical because of the very limited cooling power available there. In this case, the architecture we have assumed puts a large heat load on the 4K stage due primarily to the static heat load down the cables from the 50K stage and the presence of a large number of low noise amplifiers. For a Bluefors XLD1000-SL dilution refrigerator using HDW wiring, the theoretical limit of the system using all the available cooling power is approximately 200 qubits. However, with engineering margin in the cooling power and the available space for microwave readout circuitry at the mixing chamber, the practical limit is approximately 140 qubits.
@@
We caution readers that the model does not capture some additional heat loads that change the results. In particular, the active heat load due to the resistance of the package that holds the processor is not considered here. Other heat loads, such as the presence of a long-life cold trap, can add further static loads to the 50K and 4K stages. As such, this model only captures a general overview, and we use it to demonstrate a system with sufficient engineering margin that could be used as a starting point for a real design. The full configuration and sources of heat would need to be modeled before the authors would have sufficient confidence to recommend any particular design. In addition, the time taken to cool to base temperature may be an important system parameter. This thermal model only considers the steady state case; a more complete model would be needed to determine the cooldown time, which depends on the mass, heat capacity, and thermal conductances of each item attached to a stage.
@@
Options to improve the thermal performance include replacing the SC-086/50-SCN-CN cable with a microwave flex cable that reduces the static load by scaling the cross-sectional area and combining several signals into one assembly, or with cables made from other low thermal conductivity metals such as brass or stainless steel. In this case, there is a trade-off between reducing the static load due to lower thermal conductivity and increasing the active load due to higher electrical resistivity. Using a superconducting cable would dramatically reduce the static and active heat loads, particularly between the 4K and the MXC stages. This option is possible at the lower fridge stages below the 4K stage, where the temperatures are below the superconducting transition temperatures of commercially available superconducting cables zotero-367. The TWPA pump termination resistor could also be relocated to the cold plate, which has a lower heat load. New designs of kinetic inductance parametric amplifiers operate at 4 K and may be able to replace the current LNA and the TWPA malnou2021. Optical transduction of the microwave output signal at the mixing chamber vanthiel2025 is another method that would allow the LNAs to be removed and replace the microwave readout chain with an optical one. Finally, control lines using optical fibers and transducers to create RF signals at the base of the fridge have also been demonstrated and could potentially replace the coax cables lecocq2021.
@@
In addition, the required volume, mass, and mounting infrastructure for typical off-the-shelf readout components such as circulators and TWPAs could potentially put an upper bound on the number of readout chains that could fit in the MXC, which in turn would put an upper bound on the number of qubits possible for such a system. Based on experience with the current hardware size, it would be challenging to accommodate more than the 24 readout lines required for the 144-qubit processor on the mixing chamber of a Bluefors XLD1000-SL.
@@ VI Conclusion
We measured the thermal conductivity and electrical resistivity of SC-086/50-SCN-CN coaxial cables from room temperature to approximately 4 K. We then used this data to create smooth functions of each respective property with temperature, using simple approximations to extend the fits beyond the experimental data down to the millikelvin regime. We then used these functions to calculate static and active heat loads for a superconducting quantum processor in a dilution fridge in realistic configurations. Finally, we applied these thermal models to a set of hypothetical model flux-tunable superconducting transmon processors of increasing size. The model processors use the tunable coupler architecture, and so each transmon requires two signal lines (an XY control and a flux bias line), while the tunable couplers require a flux bias line. We also modeled the readout system, including TWPAs at the mixing chamber and semiconductor LNAs at the 4K stage. For a Bluefors XLD1000-SL dilution refrigerator using HDW wiring, the theoretical limit of the system using all the available cooling power is approximately 200 qubits. However, with engineering margin in the cooling power and the available space for microwave readout circuitry at the mixing chamber, the practical limit is approximately 140 qubits.
VII List of Abbreviations
TWPA, Traveling Wave Parametric Amplifier
LNA, Low Noise Amplifier
MXC, Mixing Chamber
CP, Cold Plate
HDW, Bluefors High-Density Wiring
PTFE, Polytetrafluoroethylene, also known as Teflon
EDX, Energy Dispersive X-ray analysis
VIII Competing Interests
The authors declare that they have no competing interests.
IX Funding
This work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359..
X Availability of Data and Materials
The thermal conductivity and electrical resistivity datasets are available at https://doi.org/10.5281/zenodo.15238913
XI Authors’ Contributions
NR, XW, and MF developed the thermal simulation code based on much earlier versions written by those in the Acknowledgments section. MF analyzed the cross-sectional imaging and EDX results. TH measured and analyzed the thermal conductivity. DS and MV measured the electrical resistivity. NR fitted the thermal conductivity and electrical resistivity data. DS, MV, and MH provided fridge data and input on the model quantum processor specifications. NR and MF ran the thermal simulation code. NR, DS, MV, and MF analyzed and interpreted the data. NR, DS, MV, and MF wrote the manuscript. All authors commented on the manuscript. All authors have read and approved the final manuscript.
XII Acknowledgements
We would like to thank Jen-Hao Yeh, Saniya Deshpande, and Michael Selvanayagam for their guidance in the early stages of this project. We would also like to thank David Gunnarsson at Bluefors for helpful comments and suggestions on the published performance data of Bluefors dilution refrigerators.